# Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

### v2

@@ 5.4 Discussion about using CLIP in GCD
Furthermore, we conduct a toy experiment to prove our TES can deal with a scenario where CLIP lacks information on a specific category class. Since CLIP saw most of the visual concepts and corresponding texts before 2022, we constructed a small dataset of new energy vehicles (NEV) that appeared in 2023. As shown in Table VIII, the NEV dataset contains 12 categories, each with 50 images from the Internet, and the classnames of the dataset consist of the brand and model of the car. We split them in the same way as standard benchmarks.
TABLE IX: The results (%) on the NEV dataset.
     CLIP zero-shot performance       10.7
     GCD performance
@@
     SimGCD      54.7      88.0      38.0
     SimGCD-CLIP      79.1      96.7      70.3
     GET(ours)      85.3      96.0      80.0
As shown in Table IX, due to the lack of information on the NEV dataset in CLIP, achieving only 10% zero-shot accuracy across all classes and CLIP can only recognize 5 out of the 12 categories. Due to the dataset’s fine-grained nature and few samples, SimGCD with the DINO backbone performs poorly on ‘New’ categories. Although fine-tuning with CLIP’s visual encoder as the backbone can yield great results, our GET with multi-modal features demonstrates a 9.7% improvement on ‘New’ categories. This proves the effectiveness of our TES in scenarios where CLIP lacks information or the CLIP text encoder has not encountered corresponding class names.
@@
Our trained TES can be considered as a special fine-tuned text encoder. This text encoder takes visual images as input and produces corresponding textual features as output. Our align loss ensures modal alignment, while the distill loss facilitates the model’s adaptation to the dataset’s distribution. For classes that CLIP has not seen before, TES can be regarded as a caption model [10]. For each input image, TES assigns a corresponding caption, expressing each caption in the form of modal-aligned text features. The text embeddings or captions corresponding to images can serve as valuable supplementary information, assisting the GCD task in a multimodal manner.
5.5 Different pre-trained models
@@
5.6 Qualitative results
Attention map visualization of GET.  We visualize the attention map of class tokens in Fig. 8. Compared to SimGCD-CLIP, our approach additionally focuses on the feather texture of birds, which is crucial for distinguishing visually similar fine-grained bird species. With the assistance of text information, the attention maps of our visual branch become more refined, focusing on more discriminative regions.
Figure 8: Attention map visualization of class tokens.
The t-SNE visualization.  We randomly sample 20 classes from the 200 classes in the CUB dataset and perform the t-SNE projections of visual and text features. As shown in Fig. 9, the features of SimGCD-CLIP are not well separated, while both visual and text features of our method exhibit clear and compacter clusters. We also provide the cluster results of GET in the Appendix.
Figure 9: The t-SNE visualization of SimGCD-CLIP and our GET on CUB dataset.
6 Conclusions
In this work, we propose to leverage multi-modal information to solve the GCD task. In particular, we introduce a text embedding synthesizer to generate pseudo text embeddings for unlabelled data. Our text embedding synthesizer module makes it possible to use CLIP’s text encoder, thus unlocking the multi-modal potential for the GCD task. Meanwhile, we use a dual-branch training strategy with a cross-modal instance consistency objective, which facilitates collaborative action and mutual learning between the different modalities, thus generating more discriminative classification prototypes. The superior performance on multiple benchmarks demonstrates the effectiveness of our method.
@@ 6 Conclusions
Broader impact.  Our approach introduces text information into the GCD task through a novel text embedding synthesizer module, extending the GCD to a multi-modal paradigm. We believe that the introduction of TES will encourage future research in solving GCD in a multi-modal manner.
7 Acknowledgments
This work is funded by NSFC (NO. 62225604, 62206135, 62202331), and the Fundamental Research Funds for the Central Universities (Nankai Universitiy, 070-63233085). Computation is supported by the Supercomputing Center of Nankai University.

### v3

@@ 8 Experimental setup
Table 9: The class names for the NEV dataset.
Implementation details.  We use a CLIP [40] pre-trained ViT-B/16 [12] as the image and text encoder. In the first stage, we train a fully connected layer to transfer image embeddings into pseudo-tokens. In the second stage, the projector of the image encoder is removed, resulting in features with a dimension of 768. The exception is ImageNet-1K, we remain and fine-tune the last projection layer, which avoids gradient explosion and improves results with lower computational cost, resulting in features with a dimension of 512. We use a single linear layer to turn pseudo text embeddings generated by TES into learnable embeddings while changing their dimensions (512 to 768) to match those of the visual features. The batch size is fixed to 128 for training and 256 for testing. Training is done with an SGD optimizer and an initial learning rate of 0.1 decayed by a cosine annealing rule. We train for 200 epochs on each dataset in both two stages. In the first stage, we set the number of pseudo text tokens to 7. The balance coefficient $\lambda$ is set to 0.35 as [45], and $\lambda_{c}$ is set to 1. The temperature value $\tau_{a}$ is set to 0.01 while other temperature values $\tau_{sc}$, $\tau_{c}$, $\tau_{s}$ , $\tau_{t}$ and the balanced value $\epsilon$ are as same as [51]. The augmentation exactly follows the previous, in which RandomCrop creates two views. All experiments are conducted with 4 NVIDIA GeForce RTX 3090 GPUs.
Method Backbone NCT-CRC-HE
All Old New
SimGCD DINO 77.1 79.9 75.1
SimGCD CLIP 79.1 93.2 69.2
GET (ours) CLIP 83.8 94.5 76.3
Table 10: Results on the medical dataset.
9 Discussion about Using CLIP in GCD
An evident fact is that using a more powerful backbone facilitates the transfer of knowledge learned from labeled data to unlabeled data [45, 51, 47, 50]. Due to the strong generalization ability of CLIP, it can encode more discriminative features, and its multi-modal information aids in discovering new categories, making it a natural choice for introducing CLIP. As discussed in the main text, using CLIP in GCD has three significances: methodological significance, forward-looking significance, and practical implications.
We argue that the key to leveraging CLIP for GCD lies in how to use its text encoder, given the presence of unlabeled data in GCD tasks. In this section, we provide supplementary analyses to complement the discussions in the main text. To be specific, we validate the effectiveness of our method and substantiate the incorporation of CLIP into GCD by addressing the following questions:
1. Does the performance gain originate from the CLIP (text encoder) being pre-exposed to the new categories?
In other words, we need to verify the effectiveness of our method on categories that are unseen by CLIP. The experimental results on the NEV and TV-100 datasets (Tab. 7 in the main paper) demonstrate the effectiveness of our method in scenarios where CLIP lacks prior information.
The intuition behind our TES can be explained from two perspectives. First, our trained TES can be considered as a special fine-tuned text encoder. This text encoder takes visual images as input and produces corresponding textual features as output. Our align loss ensures modal alignment, while the distill loss facilitates the model’s adaptation to the dataset’s distribution. Second, TES can be regarded as a caption model [35]. For each input image, TES assigns a corresponding caption, expressing each caption in the form of modal-aligned text features. The text embeddings or captions corresponding to images can serve as valuable supplementary information, assisting the GCD task in a multimodal manner.
2. In more realistic scenarios where class names (either old or new) cannot be generated or retrieved, does our method remain effective?
To address this concern, we first conducted GCD experiments on a medical dataset: the NCT-CRC-HE dataset. The NCT-CRC-HE [22] dataset comprises histological images of human colorectal cancer, containing nine categories. We selected the first five categories as the old classes. For the medical dataset, generating or retrieving new class names is challenging, its class names need expert knowledge. Our method directly generates text features aligned with visual features, and the experimental results in Tab. 10 demonstrate its effectiveness.
Moreover, in certain scenarios, it is also difficult to obtain the class names of base categories. For example, in ultra-fine-grained datasets [31, 54], different categories represent different types of soybean leaves. In such cases, category discovery using CLIP becomes significantly more challenging. To address this, we remove the distillation loss in TES, allowing the use of CLIP’s text encoder even when base class names are unavailable. Tab. 11 presents the experimental results on three ultra-fine-grained datasets, demonstrating the effectiveness of our approach. Meanwhile, the results in Tab. 5 of the main paper further demonstrate the effectiveness of our TES in scenarios where the class names of base categories are unavailable, by removing the distillation loss.
SoyAgeing-R1 SoyAgeing-R3 SoyAgeing-R4
Method All Old New All Old New All Old New
SimGCD 37.3 48.4 31.7 32.7 47.2 25.5 35.4 46.4 29.9
GET (ours) 47.9 56.5 43.6 46.0 55.2 41.4 46.6 52.8 43.4
Table 11: Results on ultra-fine-grained datasets using CLIP backbone.
10 Additional Experiments and Analysis
The architecture of TES.  In TES, we use a single linear layer to transform the visual embedding to pseudo tokens and set the number of pseudo text tokens to 7 across all datasets. Experiment results on the architecture of TES for the CUB dataset are presented in Fig. 5, proving that a single linear layer can effectively transfer visual features into text tokens while reducing the computational cost.
Figure 5: Experiments on the pseudo-tokens and layers in TES.
Effectiveness of text embedding synthesizer.  In order to prove that our text embedding synthesizer can generate reliable and discriminative representations, we visualize the text embeddings of CIFAR10 with t-SNE. As shown in Fig. 6, the initial text embeddings within the same class exhibit clear clustering, and the learnable embeddings further produce compacter clusters. Moreover, we introduce TES into the non-parametric GCD by straightforwardly concatenating text and image features before semi-supervised k-means classification. As in Tab. 13, with the help of text information, GCD gains about 5% average improvement on ‘All’ classes over 3 datasets, demonstrating the importance of multi-modal information in GCD task and our TES can be widely used in multiple GCD methods.
CUB Stanford Cars FGVC-Aircraft CIFAR10 CIFAR100 ImageNet-100
Methods All Old New All Old New All Old New All Old New All Old New All Old New
PromptCAL-CLIP 65.5 68.7 63.9 74.0 80.8 70.8 54.5 61.8 51.0 88.7 96.5 84.8 80.5 82.4 76.8 87.4 93.6 84.3
Table 12: Results of PromptCAL-CLIP.
FGVC-Aircraft ImageNet-100 ImageNet-R
Method All Old New All Old New All Old New
GCD-CLIP 45.3 44.4 45.8 75.8 87.3 70.0 44.3 79.0 25.8
+TES 49.6 49.3 49.8 80.0 95.1 72.4 49.4 79.4 33.5
Table 13: Effectiveness of TES in non-parametric GCD.
Figure 7: Different ViT finetune strategies.
Additional baseline results.  As shown in Tab. 12. We provide results of PromptCAL-CLIP on three fine-grained datasets and three image classification generic datasets. For three fine-grained datasets, our method outperforms PromptCAL-CLIP on all datasets and classes. In particular, we surpass PromptCAL-CLIP by 11.5%, 4.5%, and 4.4% on ‘All’ classes of CUB, Stanford Cars, and Aircraft, respectively. As for the generic datasets, our method surpasses PromptCAL-CLIP on all datasets and achieves the best results on CIFAR-100 and ImageNet-100 datasets.
Error bars for main results.  The experimental results presented in the paper are the averages of three independent repeated runs. We provide the performance standard deviation of our main results on all evaluation datasets with three runs in Tab. 14.
Dataset All Old New
CIFAR10 97.2$\pm$0.1 94.6$\pm$0.1 98.5$\pm$0.1
CIFAR100 82.1$\pm$0.4 85.5$\pm$0.5 75.5$\pm$0.5
ImageNet-100 91.7$\pm$0.3 95.7$\pm$0.0 89.7$\pm$0.4
CUB 77.0$\pm$0.5 78.1$\pm$1.6 76.4$\pm$1.2
Stanford Cars 78.5$\pm$1.3 86.8$\pm$1.5 74.5$\pm$2.2
FGVC-Aircraft 58.9$\pm$1.2 59.6$\pm$0.6 58.5$\pm$1.8
Herbarium 19 49.7$\pm$0.4 64.5$\pm$0.8 41.7$\pm$0.8
ImageNet-1K 62.4$\pm$0.0 74.0$\pm$0.2 56.6$\pm$0.1
ImageNet-R 58.1$\pm$2.4 78.8$\pm$0.5 47.0$\pm$3.9
Table 14: The standard deviation of our method.
Results of two branches.  We report the results of visual and text branches for ‘All’ classes across six datasets in Tab. 15. For 2 generic datasets (CIFAR10 and ImageNet-100), though the text branch does not achieve state-of-the-art performance, it still exhibits great performance. For 2 fine-grained datasets (CUB and Stanford Cars), both visual and text branches outperform previous methods by a large margin, while the visual branch performs better. For 2 challenging datasets (ImageNet-1K and ImageNet-R), both visual and text branches achieve remarkable results. Due to the challenging datasets comprising a significant number of unknown classes (ImageNet-1k dataset) or diverse visual concepts within the same class (ImageNet-R dataset), the consistency in text information for the same class contributes to the potentially higher discriminative power of the text branch, leading to better performance.
Dataset Visual Branch Text Branch
CIFAR10 97.2$\pm$0.1 95.1$\pm$0.0
ImageNet-100 91.7$\pm$0.3 90.1$\pm$0.1
CUB 77.0$\pm$0.5 73.6$\pm$0.8
Stanford Cars 78.5$\pm$1.3 73.1$\pm$0.6
ImageNet-1K 62.4$\pm$0.0 63.5$\pm$0.1
ImageNet-R 58.1$\pm$2.4 62.6$\pm$0.9
Table 15: The results of two branches.
We also provide the performance evolution of two branches throughout the model learning process on the CUB dataset (see in  Fig. 8), the mutual promotion and fusion of the two branches resulted in excellent outcomes. In our experiments, we consistently and simply select the results from the visual branch.
Figure 8: Performance evolution of two branches throughout the model learning process on the CUB dataset.
Method CIFAR10 CIFAR100 ImageNet-100 CUB Stanford Cars
Ground truth 10 100 100 200 196
GCD-CLIP 5 (50%) 94 (6%) 116 (16%) 212 (12%) 234 (19%)
$+$TES 8 (20%) 97 (3%) 109 (9%) 212 (12%) 220 (12%)
Table 16: Estimation of class number in unlabelled data. The table shows the estimated number and the error.
Results with the estimated number of classes.  Vaze et al.  [45] provides an off-the-shelf method to estimate the class number in unlabelled data. We introduce text embeddings generated by our TES into the off-the-shelf method by simply concatenating text and image features before class number Estimation. As shown in Tab. 16, multi-modal features can estimate a more accurate class, demonstrating our multi-modal method is effective in category number estimation. Following previous works, we assume the number of classes for each dataset is known and provide experimental results in the main paper. Tab. 17 shows the results using the estimated number of classes for CUB and SCars datasets.
Method Known $C$ CUB Stanford Cars
All Old New All Old New
GET 77.0 78.1 76.4 78.5 86.8 74.5
GET ✗ (w/ Est.) 75.6 75.9 75.5 76.8 87.6 71.6
Table 17: Results with the estimated number of classes
Computation complexity analysis.Tab. 18 shows the computation complexity. Our TES uses a frozen visual encoder and stage 2 finetunes the last block in another visual encoder, thus the 2 stages share the same visual encoder for the first 11 blocks but a different last block, resulting in a low computational complexity increase.
Inference Time Learnable Params FLOPs
Methods (s/per img) (M) (G)
SimGCD-CLIP $\text{5.2}\times\text{10}^{-\text{3}}$ 13.4 35.2
GET(ours) $\text{5.2}\times\text{10}^{-\text{3}}$ 15.6 38.6
Table 18: Computation complexity analysis.
The anchor prototypes.  In the CICO, the anchor prototypes are calculated by averaging the features of labeled anchor samples, making them more dynamic compared to directly using the prototype classifier $\eta$. The ablation on CUB is shown in Tab. 19.
@@ 10 Additional Experiments and Analysis
use classifier $\eta$ 76.3 77.6 75.6
use anchors (ours) 77.0 78.1 76.4
Table 19: The anchor prototypes.
Methods All Old New
SimGCD-CLIP 83.1$\pm$7.4 99.2$\pm$0.3 75.1$\pm$10.9
GET(ours) 90.0$\pm$1.9 99.2$\pm$0.2 85.5$\pm$2.8
Table 20: The results on Clevr-4 (Texture) in 5 runs.
Experiments on the Clevr-4 dataset  Recently, [48] presented a synthetic dataset Clevr-4 to examine whether the GCD method can extrapolate the taxonomy specified by the labeled set. Most attributes of Clevr-4, such as shape, color, and count, are easily clustered (achieving close to 99% accuracy with CLIP). However, texture attributes pose a certain level of challenge. Therefore, we evaluate our method on the texture attributes of Clevr-4. As shown in Tab. 20, our method achieves higher accuracy and lower standard deviation compared to SimGCD-CLIP, proving that the GCD method can cluster data at specified levels based on the constraint of labeled text information, which is worthy of attention and exploration.
CUB Stanford Cars FGVC-Aircraft CIFAR10 CIFAR100 ImageNet-100
WordNet [36] 41.8 35.2 45.1 26.5 21.9 29.0 16.5 13.3 18.2 18.0 18.6 17.8 18.6 18.9 18.8 28.8 39.1 23.6
CC3M [42] 20.8 20.8 20.9 18.7 19.0 18.5 15.4 11.6 17.3 8.1 8.3 8.1 13.0 13.6 11.6 8.9 12.1 7.2
Table 21: Results (%) of retrieval-based approach.
Retrieval baselines.  To address the challenge of missing class names, another method might involve utilizing a knowledge base of potential class names (nouns) and then using CLIP to retrieve names from this corpus. Images that share the same retrieved name could be grouped together, and clustering accuracy could then be measured based on these groupings. Therefore, this retrieval-based approach serves as an important baseline. Tab. 21 shows the results of retrieval-based approach, using WordNet [36] and CC3M [42] as corpus.
The impact of hyper-parameter $\lambda_{c}$.  In our method, we set $\lambda_{c}$ to 1 for all datasets to prevent over-tuning. Tab. 22 shows the ablation of the impact of $\lambda_{c}$.
$\lambda_{c}$ CUB Stanford Cars
All Old New All Old New
0.5 76.3 74.7 77.1 79.0 88.7 74.3
1 77.0 78.1 76.4 78.5 86.8 74.5
1.5 75.3 75.9 75.0 79.6 90.7 74.2
2 75.0 77.3 73.8 79.0 86.8 75.3
Table 22: The impact of hyper-parameter $\lambda_{c}$
11 Pseudo-code
Output: Predicted label $\hat{y}_{i}$.
/* Stage 1: TES Training */
repeat
       for  $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\in\text{each batch}$ do
            $\boldsymbol{z}_{i}^{v}=E_{i}(\boldsymbol{x}_{i})$ // visual embedding
            $\boldsymbol{t}_{i}=l(\boldsymbol{z}_{i}^{v}))$ // pseudo text tokens
            $\mathcal{L}_{align}\leftarrow$ Eq. (5) and Eq. (6)
            $\mathcal{L}_{distill\leftarrow}$ Eq. (7)
            $\mathcal{L}_{TES}=\mathcal{L}_{align}+\mathcal{L}_{distill}$
            Back-propagation and optimize $\theta_{\text{t}}$.
until reaching max epochs;
/* Stage 2: Dual-branch training */
repeat
       for  $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\in\text{each batch}$ do
             /* Visual-branch */
            Compute $\mathcal{L}_{ucon}^{v}$ and $\mathcal{L}_{scon}^{v}$ by replacing $\boldsymbol{h}$ in Eq. (1) and Eq. (2) with $\boldsymbol{h}^{v}$
            $\mathcal{L}_{\text{rep}}^{v}\leftarrow$ Eq. (8)
            Compute $\mathcal{L}_{\text{cls}}^{v}$ by replacing $\boldsymbol{p}$ in Eq. (3) and Eq. (4) with $\boldsymbol{p}^{v}$
            $\mathcal{L}_{\text{db}}^{v}\leftarrow$Eq. (9)
            /* text-branch */
             $\boldsymbol{\hat{z}}_{i}^{t}=E_{t}(l(E_{i}(\boldsymbol{x}_{i})))$
            $\mathcal{L}_{\text{db}}^{t}=\mathcal{L}_{\text{rep}}^{t}+\mathcal{L}_{\text{cls}}^{t}$
            Compute the multi-modal mean entropy regularization $H_{mm}$
            /* CICO */
             Calculate the visual and text anchors $\mathcal{P}_{v},\mathcal{P}_{t}$
            Compute the instance relationships by Eq. (10)
            $\mathcal{L}_{\text{CICO}}\leftarrow$ Eq. (11)
            $\mathcal{L}_{\text{Dual}}\leftarrow$ Eq. (12)
until reaching max epochs;
return $\hat{y}_{i}=\eta(f_{v}(\boldsymbol{x}_{i}))$.
Algorithm 1 Pseudocode for  GET.
12 Different pre-trained models
@@
Figure 11: Cluster accuracy of SimGCD-CLIP and our GET on some visually similar classes in CUB datasets. GCD methods relying solely on a single visual modality result in empty clusters(highlighted by red boxes); Our multi-modal approach GET avoids empty clusters and achieves higher classification accuracy.
13 Cluster results of GET
As shown in Fig. 11, we present the comparative cluster accuracy between our multi-modal approach and previous single-modal methods on some visually similar classes in CUB datasets. It is worth noting that relying solely on visual information, even with a powerful CLIP backbone, the previous method (SimGCD-CLIP) still struggles to differentiate some categories, resulting in empty clusters. However, leveraging the rich and discriminative text information of categories, our GET achieves more accurate classification results on CUB without any empty clusters across all categories, demonstrating the importance of multi-modal information in the GCD task. Furthermore, we showcase the clustering results of SimGCD-CLIP (see in Fig. 13) and our GET  (see in Fig. 13) for the 170th class, “Mourning Warbler”, in the CUB dataset. SimGCD-CLIP relies solely on visual information to categorize birds based on shape and posture, the model categorizes many visually similar samples as “Mourning Warbler”. Our approach, by incorporating text information, enhances the model’s discriminative ability and correctly identifies all instances of the “Mourning Warbler” class, achieving 100% classification accuracy for this visually challenging category.
Figure 12: SimGCD-CLIP cluster results visualization for class “Mourning Warbler” in CUB dataset. SimGCD-CLIP categorizes birds based on shape and posture and incorrectly identifies many visually similar categories, resulting in a clustering accuracy of 6.7% for class “Mourning Warbler”.
Figure 13: Cluster results visualization for class “Mourning Warbler” in CUB dataset of our GET. Our method uses multi-model information, achieving 100% classification accuracy for this visually challenging category.
14 Limitations and Broader Impact
@@
Broader impact.  Our approach introduces text information into the GCD task through a novel text embedding synthesizer module, extending the GCD to a multi-modal paradigm without extra corpus or models. We believe that the introduction of TES will encourage future research in solving GCD in a multi-modal manner.